PREDISSOCIATION DYNAMICS OF OH A^{2}Σ^{+} (v = 4)

\underline{ERIKA L. DERRO}, LOGAN P. DEMPSEY, ILANA B. POLLACK, MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

Predissociation of the excited A^{2}Σ^{+} electronic state of the OH radical results from curve crossings with repulsive potentials, namely ^{4}Σ^{−}, ^{2}Σ^{−}, ^{4}Π, producing O(^{3}P_{J}) + H(^{2}S) fragments. Previous experimental and theoretical studies of the v = 4 level of the OH A^{2}Σ^{+} state indicate that predissociation occurs at least 10-fold faster than in lower vibrational levels, resulting in a fluorescence quantum yield that is essentially zero. In the present study, Fluorescence Depletion Infrared (FDIR) spectroscopy, a UV-IR double resonance technique, is employed to characterize the linewidths and corresponding lifetimes of highly predissociative rovibrational levels of the excited A^{2}Σ^{+} electronic state of the OH radical. A least-squares fit to a Voigt line profile is used to extract the Lorentzian linewidths of the individual dips in the FDIR spectra. The homogeneous linewidth measurements, ranging from 0.23 to 0.31 cm^{−1} FWHM, demonstrate that the N = 0 − 7 rotational levels of the OH A^{2}Σ^{+} (v = 4) state undergo rapid predissociation, with lifetimes ≤ 23 ps. The experimental linewidths are in near quantitative agreement with first principle theoretical predictions.